# Nagaoka Ferromagnetism in $3\times 3$ Arrays and Beyond

### v2

@@ Abstract
Nagaoka ferromagnetism (NF) is a long-predicted example of itinerant ferromagnetism (IF) in the Hubbard model that has been studied theoretically for many years. The condition for NF, an infinite on-site Coulomb repulsion and a single hole in a half-filled band, does not arise naturally in materials. NF was only realized recently for the first time in experiments on a $2\times 2$ array of gated quantum dots. Dopant arrays and gated quantum dots in Si allow for engineering controllable systems with complex geometries. This makes dopant and quantum dot arrays good candidates to study NF in different array geometries through analog quantum simulation. Here we present theoretical simulations done for $3\times 3$ arrays and larger $N\times N$ arrays, and predict the emergence of different forms of ferromagnetism in different geometries. We find NF in perfect $3\times 3$ arrays, as well as in $N\times N$ arrays for one hole doping of a half-filled band. The ratio of the hopping $t$ to Hubbard on-site repulsion $U$ that defines the onset of NF scales as $1/N^{3.5}$ as $N$ increases, approaching the bulk limit of infinite $U$ for large $N$. Additional simulations are done for geometries made by removing sites from $N\times N$ arrays. Different forms of ferromagnetism are found for different geometries. Loops show ferromagnetism, but only for three electrons. For loops, the critical $t/U$ for the onset of ferromagnetism scales as $N$ as the loop length increases. We show that the different dependences on size for loops and $N\times N$ arrays can be understood by scaling arguments that highlight the different energy contributions to each different form of ferromagnetism. Our results show how analog quantum simulation with small arrays can elucidate the role of effects including wavefunction connectivity; system geometry, size and symmetry; bulk and edge sites; and kinetic energy in determining quantum magnetism of small systems.
@@ I Introduction
Following Nagaoka’s original work, many have tried to generalize Nagaoka’s result. Tasaki provided a proof of a generalized version of Nagaoka’s theorem [2]. Mielke and others found a new class of ferromagnetic states in the Hubbard model which they called “flat-band ferromagnetism” [5]. The possibility of extending NF to finite $U$ with finite densities of holes has also been studied for different infinite lattices [6, 7]. There have also been many theoretical works done on the instability of NF, that show when NF does not take place [8, 9, 10]. Recently, additional theoretical work has elucidated the polaronic correlation around holes in NF [11, 12].
@@ III.1 $3\times 3$ case
The $2\times 2$ array has no internal sites. However, the $3\times 3$ dopant array studied by [21] has one internal site. In this section, we look at $3\times 3$ and larger arrays to see how NF behaves as the size of the system increases and the system approaches the 2D bulk limit. NF has not yet been presented for the largest finite systems discussed here. The largest arrays consider here are $8\times 8$ arrays, which should be accessible with dopant arrays. These arrays have more internal sites than edge sites and should represent to transition to the 2D bulk limit. Our results will define the $t/U$ ratios needed for NF. For the $3\times 3$ array, the ground state energy for each spin-$z$ subspace is calculated to determine the presence of NF, and if it exists, to identify the point at which the transition occurs. The half-filled band with one hole has $8$ electrons, and the maximum total spin is $4$. If all $9s_{z}$ subspaces have the same ground-state energy, the system has NF.
@@ III.2 $N\times N$ arrays
Figure 5: NF for $N\times N$ arrays. (a) NF for $4\times 4$ and $5\times 5$ arrays. Energy of GS for each $s_{z}$ subspace plotted as a function of $|t/U|$, offset by the energy of the maximum $s_{z}$ state. The black dashed line indicates the transition for NF. Both system are in the one hole case, $4\times 4$ arrays have $S_{max}^{4\mathrm{x}4}=15/2$, and for $5\times 5$ arrays $S_{max}^{5\mathrm{x}5}=12$. Here $s_{z}$ subspaces accessible with ED are shown as solid lines, DMRG as dashed lines. (b) Transition point v.s. $N$, including both ED and DMRG results. Fitted power law functions shows how the system approaches the bulk limit (Nagaoka limit $U\rightarrow\infty$) as $N$ gets larger.
Simulation done for $4\times 4$ and $5\times 5$ arrays for one-hole in a half-filled band are shown in Fig.5a. The larger $N\times N$ arrays behave similarly to the $3\times 3$ and $2\times 2$ arrays, with the transition point decreasing as N gets larger, as shown in Fig. 5b. As the $N\times N$ array gets larger, the on-site interaction $U$ needs to be larger for NF to occur. This agrees with Nagaoka’s original theorem, that ferromagnetism in an infinite lattice occurs for infinite on-site interaction $U$. Fig. 5a includes the results from ED and density matrix renormalization group (DMRG) simulations. For smaller $s_{z}$ the large number of possible spin configurations makes ED difficult even for $4\times 4$ and $5\times 5$ arrays. When ED calculations are not possible, we perform DMRG simulations. Here, we flatten the 2D arrays into 1D chains with long-range hoppings, and obtain their ground state energies through the well-established two-site iterative variational searches [27, 28]. The DMRG calculations confirm the prediction that for larger $N\times N$ arrays, NF exists for a smaller region of $|t/U|$. The dependence of the transition ratio on array size $N$ shown in Fig. 5b out to the $8\times 8$ array has been well fitted to an inverse power law on $N$. In the Appendix A, a scaling argument is used to support an inverse power law dependence on $N$. The $7\times 7$ is the first lattice with more internal sites than edge sites. Our results suggest that the bulk limit for NF ferromagnetism with $U$ becoming very large is being approached by $N=8$. Table 1 summarizes the results for $N\times N$ arrays.
@@
$7\times 7$ [Uncaptioned image] yes 48 $0.00055$ DMRG
$8\times 8$ [Uncaptioned image] yes 63 $0.00023$ DMRG
@@ IV.1.3 Role of Wavefunction Connectivity
Figure 7: (a) An example showing when the connectivity condition for NF is violated. To exchange the two spins (in the box), a series of nearest neighbor hoppings is not sufficient, as the spin up and spin down electrons must be on the same site which is strongly suppressed by large repulsion $U$. (b) Example of the connectivity condition being satisfied for 3 electrons on a loop. For any sized loop, 3 electrons is the only filling where the wavefunction is connected. Any pair of spins (shown in the box) can be exchanged when all electrons are hopped around the loop.
@@ IV.1.4 Removing Two Opposite Corners
If the two sites are removed from opposite corners of the $3\times 3$ array, the structure becomes two $2\times 2$ plaquettes connected by sharing a common corner site. The ground state for this structure is ferromagnetic only when there are two holes in a half-filled band of the array, that is, when there are five electrons. The transition is at $|t/U|\approx 0.0127$ and calculations of the pair correlation function show ferromagnetic behaviour below the transition. We refer to this as the itinerant ferromagnetism (IF) case because it is not possible for one-hole doping of a half-filled band as in NF. This ferromagnetism for two-hole doping in a half-filled band can be extended to longer chains of squares. The ground state of the longer chains of corner-sharing squares is ferromagnetic but only when there are two holes in a half-filled band, and not ferromagnetic for the one-hole case (the NF state). The many-body wavefunction is not connected when there is one-hole in a half-filled band, preventing GS saturated ferromagnetism. The site shared by the two $2\times 2$ plaquettes acts as a bottleneck. A hole on one of the $2\times 2$ plaquettes can’t be used to move a spin around another plaquette that has spins at each site. A second hole is needed. Hence ferromagnetism is possible for two holes, as we have found. As before, this is a necessary but not sufficient condition. Saturated ferromagnetism does not exist for more than two holes. As the chain gets longer, the transition point also decreases as happens when arrays increase in size as seen in Fig 8. (For details, see Appendix B.2).
Figure 8: Two-hole ferromagnetism in a chain with $N$ $2\times 2$ plaquettes sharing corners on a common diagonal. Transition point versus the number $N$ of plaquettes. The fitted power law function is shown as a dotted line.
@@ IV.3.1 Three-electron Loop Ferromagnetism
When the center site is removed from the $3\times 3$ array, the array becomes an eight-site loop. The many-body wavefunction for this loop is not connected for one-hole in a half-filled band, as shown by Fig. 7b. As a consequence, NF can’t exist. Our calculations confirm this. Interestingly, the 8-site loop does have a ferromagnetic ground state for filling by 3 electrons. This result exists regardless of the size of the loop, and the transition point increases linearly as the number of sites on the loop increases, shown in Fig. 9. In fact, the many-body wavefunction on the loop is connected for a filling of three electrons. The many-body wavefunction is not connected for any higher electron filling. Thus, a ferromagnetic GS should only be possible for a filling of 3 electrons or less.
The specific cases for pentagons (five-site loops) and hexagons (six-site loops) were discussed in [16], which described ferromagnetism appearing in small pentagonal and hexagonal plaquettes at filling factors of roughly 3/10 and 1/4. However, this density dependence appears to be an artifact due to the different geometries considered. In each case, the ferromagnetism occurs for 3 electron filling. This is not a density effect. This filling for ferromagnetism arises because three-electron filling is the only filling on loops that allows a connected many-body wavefunction.
@@
Figure 9: Dependence of the transition point $|t/U|$ on loop size. As the loop gets larger, the transition point gets larger. Here only loop sizes accessible with ED are shown. A linear fit is shown.
@@ IV.3.2 Scaling Arguments for the Transition to Ferromagnetism
The arguments to support the $1/N^{3.5}$ size-scaling of the transition $t/U$ for NF in $N\times N$ arrays with one hole in half-filled band are more complicated. Here we summarize the key results. The details are given in Appendix A. The arguments are easiest to develop in the infinite $U/t$ limit. The transition is identified by why and when the excited, non-ferromagnetic states cross the ferromagnetic ground states as $U$ decreases. In the infinite $U$ limit, the energy levels of states with one hole in a half-filled band can be determined by lowest order degenerate perturbation theory where $|t/U|$ is the small perturbation parameter. In the infinite $U$ limit, the fully polarized, single-hole state is the ground state with the energy for one hole hopping freely in the array. Low-lying partially polarized excited states are shifted up from the ground state by energies that scale between $|t|/N^{3}$ and $|t|/N^{4}$ due to how the hole is localized by scattering from minority spins in the array. The fully polarized ground state in the infinite $U$ limit has no repulsive interaction energy so its energy remains fixed as $U$ becomes finite. However, the higher lying, partially polarized states gain exchange energy $-t^{2}/U$ that lowers their energy, because hopping to double occupied sites becomes possible for finite $U$. The transition to a non-ferromagnetic state occurs when the energy gain due to exchange lowers the excited state energies enough to cross the fully polarized level. This occurs at a $|t|/U$ that scales between $|t|/N^{3}$ and $|t|/N^{4}$. This is consistent with the $1/N^{3.5}$ scaling obtained from the computational results.
@@ V Summary and Discussion
In this work we have presented theoretical simulations done to probe Nagaoka ferromagnetism in arrays with different geometries. Starting from the $2\times 2$ plaquette, which was experimentally realized in [4], we have shown that perfect $3\times 3$ and $N\times N$ arrays up to $8\times 8$ exhibit NF for one-hole in a half-filled band, with a $t/U$ ratio for the onset of NF that decreases as the array size N increases, with the transition scaling as $1/n^{3.5}$. So far, 2D dopant arrays have only been studied for the $3\times 3$ configuration. However, recent advances in fabrication of dopant structures give significant promise for making highly engineered dopant arrays that will be an excellent testbed as an analog quantum simulator for this quantum magnetism .[25] Extending coupled quantum dot systems to large, complex and extended arrays shows similar promise.
@@
Electron filling plays an essential role in determining the type of ferromagnetism that can occur and the scaling of the transition $|t/U|$ that marks the crossover to ferromagnetism. Loop ferromagnetism occurs for the same, fixed, small electron filling for all loop sizes $N$. The ferromagnetic state occurs for three electron filling but not for more electrons. Nagaoka ferromagnetism in $N\times N$ occurs when there is one hole in a half-filled band. As one consequence the scaling of the transition $|t/U|$ is linear in $N$ for loops and $1/N^{3.5}$ in arrays. Loop ferromagnetism occurs for a broader range of $U$ as $N$ increases, while for NF, the ferromagnetic state occurs for a smaller range of $U$ as $N$ increases. Scaling arguments support and elucidate the scaling obtained from the computational studies. Three-electron loop ferromagnetism occurs when the kinetic energy costs of becoming ferromagnetic is compensated by the elimination of repulsive interaction in the ferromagnetic state. For NF, the transition out of the ferromagnetic state occurs when the hopping energy cost of the hole, partially localized by scattering from the minority spins is compensated by the exchange energy that is gained when hopping to create doubly occupied states is allowed at finite $U$.

### v3

@@ Abstract
Nagaoka ferromagnetism (NF) is a long-predicted example of itinerant ferromagnetism (IF) in the Hubbard model that has been studied theoretically for many years. The condition for NF, an infinite on-site Coulomb repulsion and a single hole in a half-filled band, does not arise naturally in materials. NF was only realized recently for the first time in experiments on a $2\times 2$ array of gated quantum dots. Dopant arrays and gated quantum dots in Si allow for engineering controllable systems with complex geometries. This makes dopant and quantum dot arrays good candidates to study NF in different array geometries through analog quantum simulation. Here we present theoretical simulations done for $3\times 3$ arrays and larger $N\times N$ arrays, and predict the emergence of different forms of ferromagnetism in different geometries. We find NF in perfect $3\times 3$ arrays, as well as in $N\times N$ arrays for one hole doping of a half-filled band. The ratio of the hopping $t$ to Hubbard on-site repulsion $U$ that defines the onset of NF scales as $1/N^{4}$ as $N$ increases, approaching the bulk limit of infinite $U$ for large $N$. Additional simulations are done for geometries made by removing sites from $N\times N$ arrays. Different forms of ferromagnetism are found for different geometries. Loops show ferromagnetism, but only for three electrons. For loops, the critical $t/U$ for the onset of ferromagnetism scales as $N$ as the loop length increases. We show that the different dependences on size for loops and $N\times N$ arrays can be understood by scaling arguments that highlight the different energy contributions to each different form of ferromagnetism. Our results show how analog quantum simulation with small arrays can elucidate the role of effects including wavefunction connectivity; system geometry, size and symmetry; bulk and edge sites; and kinetic energy in determining quantum magnetism of small systems.
@@ I Introduction
Following Nagaoka’s original work, many have tried to generalize Nagaoka’s result. Tasaki provided a proof of a generalized version of Nagaoka’s theorem [2]. Mielke and others found a new class of ferromagnetic states in the Hubbard model which they called ”flat-band ferromagnetism” [5]. The possibility of extending NF to finite $U$ with finite densities of holes has also been studied for different infinite lattices [6, 7], for large 2D lattices [8], and for square lattices [9]. also see [10] for Nagaoka polaron regime for finite hole density. There have also been many theoretical works done on the instability of NF, that show when NF does not take place [11, 12, 13]. Additional theoretical work has elucidated the polaronic correlation around holes in NF [10, 14, 15].
@@ III.1 $3\times 3$ case
The $2\times 2$ array has no internal sites. However, the $3\times 3$ dopant array studied by [24] has one internal site. In this section, we look at $3\times 3$ and larger arrays to see how NF behaves as the size of the system increases and the system approaches the 2D bulk limit. The largest arrays consider here are $8\times 8$ arrays, which should be accessible with dopant arrays. These arrays have more internal sites than edge sites and should represent the beginning of transition to the 2D bulk limit. Previous work [30] has been done on $9\times 9$ systems for a t-J model in the polaron picture.
Our results will define the $t/U$ ratios needed for NF. For the $3\times 3$ array, the ground state energy for each spin-$z$ subspace is calculated to determine the presence of NF, and if it exists, to identify the point at which the transition occurs. The half-filled band with one hole has $8$ electrons, and the maximum total spin is $4$. If all $9s_{z}$ subspaces have the same ground-state energy, the system has NF.
@@ III.2 $N\times N$ arrays
Figure 5: NF for $N\times N$ arrays. (a) NF for $4\times 4$ and $5\times 5$ arrays. Energy of GS for each $s_{z}$ subspace plotted as a function of $|t/U|$, offset by the energy of the maximum $s_{z}$ state. The black dashed line indicates the transition for NF. Both system are in the one hole case, $4\times 4$ arrays have $S_{max}^{4\mathrm{x}4}=15/2$, and for $5\times 5$ arrays $S_{max}^{5\mathrm{x}5}=12$. Here $s_{z}$ subspaces accessible with ED are shown as solid lines, DMRG as dashed lines. (b) Transition point v.s. $N$, including both ED and DMRG results. Fitted power law functions (discarding lower values of N) shows how the system approaches the bulk limit (Nagaoka limit $U\rightarrow\infty$) as $N$ gets larger. Inset axis shows the plot in log-log scale.
Simulation done for $4\times 4$ and $5\times 5$ arrays for one-hole in a half-filled band are shown in Fig.5a. The larger $N\times N$ arrays behave similarly to the $3\times 3$ and $2\times 2$ arrays, with the transition point decreasing as N gets larger, as shown in Fig. 5b. As the $N\times N$ array gets larger, the on-site interaction $U$ needs to be larger for NF to occur. This agrees with Nagaoka’s original theorem, that ferromagnetism in an infinite lattice occurs for infinite on-site interaction $U$. Fig. 5a includes the results from ED and density matrix renormalization group (DMRG) simulations. For smaller $s_{z}$ the large number of possible spin configurations makes ED difficult even for $4\times 4$ and $5\times 5$ arrays. When ED calculations are not possible, we perform DMRG simulations. Here, we flatten the 2D arrays into 1D chains with long-range hoppings, and obtain their ground state energies through the well-established two-site iterative variational searches [31, 32]. The DMRG calculations confirm the prediction that for larger $N\times N$ arrays, NF exists for a smaller region of $|t/U|$. The dependence of the transition ratio on array size $N$ shown in Fig. 5b out to the $8\times 8$ array has been well fitted to an inverse power law on $N$. In the Appendix A, a scaling argument is used to support an inverse power law dependence on $N$. The $7\times 7$ is the first lattice with more internal sites than edge sites. Our results suggest that the bulk limit for NF ferromagnetism with $U$ becoming very large is being approached by $N=8$. Table 1 summarizes the results for $N\times N$ arrays. Previous work on the $9\times 9$ array with open boundary conditions in [30] suggests that the transition point to a ferromagnetic ground state on the whole system is $J/t=4t/U\approx 0.001$, which is slightly larger than the transition point we predict here from scaling from the $8\times 8$ array to the $9\times 9$ array.
@@
$7\times 7$ [Uncaptioned image] yes 48 $0.0006$ DMRG
$8\times 8$ [Uncaptioned image] yes 63 $0.0002$ DMRG
@@ IV.1.3 Role of Wavefunction Connectivity
Figure 7: (a) An example showing when the connectivity condition for NF is violated. To exchange the two spins (in the box), a series of nearest neighbor hoppings is not sufficient, as the spin up and spin down electrons must be on the same site which is strongly suppressed by large repulsion $U$. (b) Example of the connectivity condition being satisfied for 3 electrons on a loop. For any sized loop, 3 electrons is the largest filling where the wavefunction is connected. Any pair of spins (shown in the box) can be exchanged when all electrons are hopped around the loop.
@@ IV.1.4 Removing Two Opposite Corners
If the two sites are removed from opposite corners of the $3\times 3$ array, the structure becomes two $2\times 2$ plaquettes connected by sharing a common corner site. The ground state for this structure is ferromagnetic only when there are two holes in a half-filled band of the array, that is, when there are five electrons. The transition is at $|t/U|\approx 0.0127$ and calculations of the pair correlation function show ferromagnetic behavior below the transition. We refer to this as the itinerant ferromagnetism (IF) case because it is not possible for one-hole doping of a half-filled band as in NF. This ferromagnetism for two-hole doping in a half-filled band can be extended to longer chains of squares. The ground state of the longer chains of corner-sharing squares is ferromagnetic but only when there are two holes in a half-filled band, and not ferromagnetic for the one-hole case (the NF state). The many-body wavefunction is not connected when there is one-hole in a half-filled band, preventing GS saturated ferromagnetism. The site shared by the two $2\times 2$ plaquettes acts as a bottleneck. A hole on one of the $2\times 2$ plaquettes can’t be used to move a spin around another plaquette that has spins at each site. A second hole is needed. Hence ferromagnetism is possible for two holes, as we have found. As before, this is a necessary but not sufficient condition. For different finite chains of squares calculated here, saturated ferromagnetism does not exist for more than two holes. Much longer chains ($N>100$) with periodic/open boundary conditions and finite hole doping densities are discussed in [33, 34]. As the chain gets longer, the transition point also decreases as happens when arrays increase in size as seen in Fig 8. (For details, see Appendix B.2).
Figure 8: Two-hole ferromagnetism in a chain with $N$ $2\times 2$ plaquettes sharing corners on a common diagonal. Transition point versus the number $N$ of plaquettes. The fitted power law function is shown as a dotted line, fitting to only larger values of N = 5, 6, 7. Inset axis shows the log-log scale.
@@ IV.3.1 Three-electron Loop Ferromagnetism
When the center site is removed from the $3\times 3$ array, the array becomes an eight-site loop. The many-body wavefunction for this loop is not connected for one-hole in a half-filled band, as shown by Fig. 7b. As a consequence, NF can’t exist. Our calculations confirm this. Interestingly, the 8-site loop does have a ferromagnetic ground state for filling by 3 electrons. For different-size loops ($N_{loop}>4$), the same, fixed-number-of-electrons (3), ferromagnetic ground state also exists and the transition point increases linearly as the number of sites on the loop increases, shown in Fig. 9. In fact, the many-body wavefunction on the loop is connected for a filling of maximum three electrons only. The many-body wavefunction is not connected for any higher electron filling. Thus, a ferromagnetic GS should only be possible for a filling of 3 electrons or less.The one-electron system is trivially polarized. Two-electrons on a loop does not have a transition point to ferromagnetism, so here, we focus on three-electrons loops. This 3-electron FM was also discussed earlier in [35] as the only stable kinetic ferromagnetism possible on a loop.
The specific cases for pentagons (five-site loops) and hexagons (six-site loops) were discussed in [19], which described ferromagnetism appearing in small pentagonal and hexagonal plaquettes at filling factors of roughly 3/10 and 1/4. However, this density dependence appears to be an artifact due to the different-size loops. For pentagon (5 sites) and hexagon (6 sites), 3-electron ferromagnetism would give electronic densities at $103/(N_{loop}\times 2)=3/$ 10 or $3/12=1/4$. In each case, the ferromagnetism occurs for 3 electron filling. This is not a density effect. This filling for ferromagnetism arises because three-electron filling on a loop allows for a connected many-body wavefunction, while all higher electron fillings do not. For more complex structures like two or three adjacent loops, special fillings at a fixed number of electrons also exist, for example, two adjacent pentagons or hexagons each have 5-electron ferromagnetism, which can be extended to larger loops, like two adjacent nonagons or even two adjacent loops with different sizes. For more discussion, see Appendix B.3.
@@
Figure 9: Dependence of the transition point $|t/U|$ on loop size $N_{loop}$ for 3-electron ferromagnetism. As the loop gets larger (while the electron number is fixed at 3) and the $t/U$ transition point to ferromagnetism gets larger. Here only loop sizes accessible with ED are shown. A linear fit is shown.
@@ IV.3.2 Scaling Arguments for the Transition to Ferromagnetism
The arguments to support the $1/N^{4}$ size-scaling of the transition $t/U$ for NF in $N\times N$ arrays with one hole in half-filled band are more complicated. Here we summarize the key results. The details are given in Appendix A. The arguments are easiest to develop in the infinite $U/t$ limit. The transition is identified by why and when the excited, non-ferromagnetic states cross the ferromagnetic ground states as $U$ decreases. In the infinite $U$ limit, the energy levels of states with one hole in a half-filled band can be determined by lowest order degenerate perturbation theory where $|t/U|$ is the small perturbation parameter. In the infinite $U$ limit, the fully polarized, single-hole state is the ground state with the energy for one hole hopping freely in the array. Low-lying partially polarized excited states are shifted up from the ground state by energies that scale between $|t|/N^{3}$ and $|t|/N^{4}$ due to how the hole is localized by scattering from minority spins in the array. The fully polarized ground state in the infinite $U$ limit has no repulsive interaction energy so its energy remains fixed as $U$ becomes finite. However, the higher lying, partially polarized states gain exchange energy $-t^{2}/U$ that lowers their energy, because hopping to double occupied sites becomes possible for finite $U$. The transition to a non-ferromagnetic state occurs when the energy gain due to exchange lowers the excited state energies enough to cross the fully polarized level. This occurs at a $|t|/U$ that scales between $|t|/N^{3}$ and $|t|/N^{4}$. This is consistent with the $1/N^{4}$ scaling obtained from the computational results.
@@ V Summary and Discussion
In this work we have presented theoretical simulations done to probe Nagaoka ferromagnetism in arrays with different geometries. Starting from the $2\times 2$ plaquette, which was experimentally realized in [4], we have shown that perfect $3\times 3$ and $N\times N$ arrays up to $8\times 8$ exhibit NF for one-hole in a half-filled band, with a $t/U$ ratio for the onset of NF that decreases as the array size N increases, with the transition scaling as $1/n^{4}$. So far, 2D dopant arrays have only been studied for the $3\times 3$ configuration. However, recent advances in fabrication of dopant structures give significant promise for making highly engineered dopant arrays that will be an excellent testbed as an analog quantum simulator for this quantum magnetism .[28] Extending coupled quantum dot systems to large, complex and extended arrays shows similar promise.
@@
Electron filling plays an essential role in determining the type of ferromagnetism that can occur and the scaling of the transition $|t/U|$ that marks the crossover to ferromagnetism. Loop ferromagnetism occurs for the same, fixed, small electron filling for all loop sizes $N$. The ferromagnetic state occurs for three electron filling but not for more electrons. Nagaoka ferromagnetism in $N\times N$ occurs when there is one hole in a half-filled band. As one consequence the scaling of the transition $|t/U|$ is linear in $N$ for loops and $1/N^{4}$ in arrays. Loop ferromagnetism occurs for a broader range of $U$ as $N$ increases, while for NF, the ferromagnetic state occurs for a smaller range of $U$ as $N$ increases. Scaling arguments support and elucidate the scaling obtained from the computational studies. Three-electron loop ferromagnetism occurs when the kinetic energy costs of becoming ferromagnetic is compensated by the elimination of repulsive interaction in the ferromagnetic state. For NF, the transition out of the ferromagnetic state occurs when the hopping energy cost of the hole, partially localized by scattering from the minority spins is compensated by the exchange energy that is gained when hopping to create doubly occupied states is allowed at finite $U$.